# Model Predictive Adversarial Imitation Learning for Planning from Observation

### v1

@@ Appendix A Infinite Horizon Model Predictive Path Integral
In this section we present the full algorithm in detail, including MPPI as described in [17]. Modifications to “conventional” MPPI for MPAIL are highlighted in blue. Where applicable, $(\mathbf{x})_{i}$ indicates the $i$th entry of $\mathbf{x}$ (in its first dimension, if $\mathbf{x}$ is a tensor).
@@
2:Number of trajectories to sample $N$;
@@
14:for $j\leftarrow 0$ to $J-1$ do
15:  for $k\leftarrow 0$ to $N-1$ do $\triangleright$ Model rollouts and costing (parallelized)
@@
17:   for $t^{\prime}\leftarrow 0$ to $H-1$ do
18:     $a^{k}_{t^{\prime}}\sim\mathcal{N}((\mathbf{a}_{t})_{t^{\prime}},\Sigma)$triangleright Sample action at predicted state
19:     $\tilde{s}^{k}_{t^{\prime}+1}\leftarrow f_{\psi}(\tilde{s}^{k}_{t^{\prime}},a^{k}_{t^{\prime}})$triangleright Predict next state
20:     $c_{t^{\prime}}^{k}\leftarrow$ $c_{\theta}(s_{t^{\prime}}^{k},s_{t^{\prime}+1}^{k})$triangleright Compute state-transition costs
@@
[…]
@@
26:  for $k\leftarrow 0$ to $N-1$ do $\triangleright$ Weight using exponential negative cost
@@
29:  for $t^{\prime}\leftarrow 0$ to $H-1$ do $\triangleright$ Optimal plan from weighted-average actions
@@
34:  $(\boldsymbol{\sigma}_{t})_{i}\leftarrow\sqrt{\sum_{k=0}^{N-1}w(\tau_{k})[((\mathbf{a}_{t}^{J})_{0})_{i}-(a^{k}_{0})_{i}]^{2}}$
@@
36:return $\mathbf{a}^{J}_{t}$, $\boldsymbol{\sigma}_{t}$
@@
9:  $s_{t}\sim\mathcal{T}(\cdot|s_{t-1},a_{t-1})$triangleright Step and perceive environment
@@
12:   $a_{t}\sim\mathcal{N}((\mathbf{a}^{*}_{t})_{0},I\boldsymbol{\sigma}_{t})$
@@
16:  $r_{t}\leftarrow r_{\theta}(s_{t},s_{t+1})$triangleright Reward from discriminator
@@ B.1 MPPI as a Policy
Let $f(s,s^{\prime})=c(s,s^{\prime})+\beta\mathbb{KL}(\pi(\cdot|s)||\overline{\pi}(\cdot|s))$. For either objective to be valid the cost and KL Divergence would have to be bounded. Thus, we may safely assume that $f$ is uniformly bounded $||f||_{\infty}\leq K$. Let $\delta_{t}=\,\mathbb{E}_{s_{t},s_{t+1}\sim d^{t}}[f(s_{t},s_{t+1})]-\mathbb{E}_{s,s^{\prime}\sim d^{\pi}}[f(s,s^{\prime})]$ be the error between estimates of the objective.
@@
$\exists\lambda\in(0,1),M\in\mathbb{N}\;\text{s.t}\;||d^{t}-d^{\pi}||_{\text{TV}}\leq M\lambda^{t}$ (24)
@@ B.2 MPAIL as an Adversarial Imitation Learning Algorithm
In this section, we integrate findings from Section B.1 with the AIL objective to theoretically validate MPAIL as an AIL algorithm. Specifically, we observe that at optimality, we recover the log expert-policy transition density ratio, which in turns yields a maximum entropy policy on state-transitions. We then discuss the identifiability limits imposed by observing only $(s,s^{\prime})$ rather than $(s,a,s^{\prime})$. Throughout this section we make use of the state-transition occupancy measure, defined as $\rho_{\pi}:\mathcal{S}\times\mathcal{S}\rightarrow\mathbb{R}$ where $\rho_{\pi}(s,s^{\prime})=\sum_{t=1}^{\infty}\gamma^{t}\mathcal{T}(S_{t+1}=s^{\prime},S_{t}=s\,|\,\pi)$ as in [23].
@@
Note that the optimal discriminator is achieved when $D^{\ast}(s,s^{\prime})=\frac{\rho_{E}(s,s^{\prime})}{\rho_{E}(s,s^{\prime})+\rho_{\pi}(s,s^{\prime})}$ as used in [42] and shown in [43] Section 4 Proposition 1.
@@
Remark B.2.3. On Identifiability. A question naturally arises about the limitations that being state-only imposes. If state transitions are deterministic and invertible, observing $(s,s^{\prime})$ is the same as observing the unique action $a$ that caused it. Then $r(s,s^{\prime})=r(s,a)$ and by the 1-1 correspondence of policies with state-action occupancy measures [12], the recovered policy becomes unique.
@@ C.1 Regularization
Spectral Normalization. As often found in GAN and AIL surveys [31, 44], we corroborate that applying spectral normalization to the discriminator architecture appeared to have improved MPAIL training stability and performance. Application of spectral normalization to the value network did not appear to make a noticeable difference.
@@ C.2 Hyperparameters
Temperature Decay. As noted in Section 3, we found that an initial temperature with a gradual decay (down to a minimum) was helpful in preventing early and unrecoverable collapse. The intuition for this decision is similar to that of decaying policy noise injection in many popular RL frameworks [45, 20], since the temperature is directly related to the variance of the optimized gaussian distribution. This component remains under investigation as its usage is not always necessary for meaningful convergence, but it is perhaps practically useful as it alleviates temperature tuning labor.
Value-to-Discriminator Update Ratio. Common to existing AIL (and GAN) implementations, MPAIL benefits from a balancing of generator and discriminator updates. Note that, like GANs, AIL tends to oscillate aggressively throughout training [46]. As MPAIL does not enforce a constrained policy update each epoch (as TRPO does [47]), the policy is exposed more directly to the discriminator’s oscillations which can further hinder on-policy value estimation. A further converged value function is also theoretically more stationary from the perspective of $\pi_{\text{MPPI}}$ as infinite-horizon MPPI.
Markup. A notable quirk discovered during implementation is the relationship between costs and rewards. While the two concepts are generally regarded as dual (with negation), it is worthwhile noting that discounting is not closed under negation. Meaning, it is not correct to apply the same discount factor to the costs as they are done to the summation of rewards in the return $G_{t}$. Consider the reward with a discount applied $r_{1}=\gamma~r(s,s^{\prime})$ and the reward of a cost with a discount applied $r_{2}=-\gamma~c(s,s^{\prime})$. Observe that for $\gamma<1$, $r_{1}$ decreases while $r_{2}$ increases. Thus, when using costs, the $H$-step factor in the MPPI horizon, $\eta$, should not decrease over $t^{\prime}$. In fact, when applying $\eta<1$, we found that MPAIL does not meaningfully converge ever on the navigation task. Geldenbott and Leung [48] names the usage of $\eta>1$ as a markup. In our case, we apply a similar empirical factor such that $\eta\coloneqq 1/\gamma>1$. While we suspect a more rigorous relationship between $\eta$ and $\gamma$, we leave its derivation for future work. However, we remark that a reward-only variant of MPPI which precludes these relationships is equally possible as done in [49]. Costs are maintained in this work due to wider familiarity in practice [14, 17, 50, 51].
@@ C.3 Computation
Figure 7: Comparing “Inference” Times for Navigation Task. Time taken to complete one episode of 100 timesteps with 64 parallel environments across varying horizon lengths and MPPI optimization iterations. PPO (in policy-based AIL) is used as implemented in the RSL library [52]. All training runs in this work are performed on an NVIDIA RTX 4090 GPU. Isaac Lab is chosen as our benchmarking and simulation environment due to its parallelization and robot learning extensions [53].
@@
        (0): Linear(in_features=24, out_features=32, bias=True)
@@
        (2): Linear(in_features=32, out_features=32, bias=True)
@@
        (4): Linear(in_features=32, out_features=1, bias=True)
@@
        (0): Linear(in_features=12, out_features=32, bias=True)
@@
        (2): Linear(in_features=32, out_features=32, bias=True)
@@
        (4): Linear(in_features=32, out_features=1, bias=True)
@@
Figure 8: PyTorch [54] Model Architecture from Train Log. MPAIL readily admits other well-studied components of the model-based planning framework (e.g. sampling, dynamics) [55, 56]. This work focuses on costing from demonstration.
MPAIL is crucially implemented to be parallelized across environments in addition to trajectory optimization. In other words, in a single environment step, each parallel environment independently performs parallelized sampling, rollouts, and costing entirely on GPU without CPU multithreading. MPPI also allows for customize-able computational budget, similar to [20] (see Figure 8). For the navigation and cartpole tasks, we find that online trajectory optimization implemented this way induces little impact on training times. In exchange, MPPI can be more space intensive due to model rollouts having space complexity of $\mathcal{O}(HN|\mathcal{S}|)$ per agent. On the navigation task benchmark settings, this is an additional 245 kB per agent or 15.7 MB in total for 64 environments. Figure 8 shows benchmarks on training times that demonstrate comparable times to PPO’s policy inference. Overall, training runs for Section 4.4 between MPAIL and GAIL on the navigation task are comparable at about 45 minutes each for 500 iterations.
@@ D.1 Real-Sim-Real Navigation
Setup Details. Before continuing with the discussion of our results, we provide further details about the setup of the experiment. The platform itself is an open-source MuSHR platform as detailed in [37]. Notably, the compute has been replaced with an NVIDIA Jetson Orin NX as mentioned in Section 4.3. Poses (position, orientation; $\begin{bmatrix}x&y&z&r&p&y\end{bmatrix}$) are provided by a motion-capture system at a rate of 20 Hz. Velocities are body-centric as estimated by onboard wheel encoders ($\mathbf{v}=v_{x}\mathbf{b}_{1}+v_{y}\mathbf{b}_{2}+v_{z}\mathbf{b}_{3}$, such that $\mathbf{b}_{1}$ points forward, $\mathbf{b}_{2}$ points left, and $\mathbf{b}_{3}$ completes the right-hand frame; basis vectors are rigidly attached to the vehicle [14]). Note that the vehicle is operated without slipping nor reversing such that $v_{y}\approx v_{z}\approx 0$ and $v_{x}>0$ [33]. The recorded states used for the expert demonstration data is 240 timesteps long. Altogether, the data can be written as $s_{E}\in\{(x_{t},y_{t},z_{t},v_{x,t},v_{y,t},v_{z,t})\}_{t=1}^{240}$.
[…]
@@
Note that the direction of travel cannot be uniquely determined by a single state $s$ due to the partially observable body-centric velocity. Rather, only with the state-transition $(s_{E},s_{E}^{\prime})$ is it possible to deduce the direction of travel. For instance, consider a simplified hand-designed cost using the partially observable expert data $c(s,s^{\prime}|s_{E},s_{E}^{\prime})$. A reference vector can be computed through the difference of positions between $s^{\prime}$ and $s$ then scaled by the demonstrated velocities: $c(s,s^{\prime}|s_{E},s_{E}^{\prime})\coloneqq\lVert^{\mathcal{I}}\mathbf{v}(s,s^{\prime})-{v_{x}}_{E}[(x_{E}^{\prime}-x_{E})\mathbf{e}_{1}+(y_{E}^{\prime}-y_{E})\mathbf{e}_{2}]\rVert_{2}$ where $\mathbf{e}_{i}$ are global basis vectors for global frame $\mathcal{I}$ and ${}^{\mathcal{I}}\mathbf{v}(s,s^{\prime})$ is the robot velocity in $\mathcal{I}$. Of course, this example assumes the ability to correctly choose the corresponding $(s_{E},s_{E}^{\prime})$ pair for input $(s,s^{\prime})$ out of the entirety of the expert dataset $d^{E}$. It should be clear that partial observability and state-transitions play critical roles in the recovery of this non-trivial cost function. This experiment presents a necessary challenge towards practical AIL [31] and scalable Learning-from-Observation (LfO) [24].
@@
[…]
Meanwhile, MPAIL’s success and IRL-MPC’s improvement over GAIL is attributed to model-based planning capabilities. If the robot should find itself away from the expert distribution, the online planner enables the agent to sample back onto the demonstration. This becomes especially important when the demonstration data is severely under-defined as in this partially observable setting, which results in positive reward signal in most states in the environment. This is further supported by the fact that discriminator predictions are comparatively uncertain when training over real data. Recall that low discriminator confidence is reflected by low magnitude logit $f_{\theta}$ (cost) predictions. For reference, real experiment cost values are in the range of (0.022,0.0180)(-0.022,-0.0180)( - 0.022 , - 0.0180 ) whereas cost values in benchmarking runs with synthetic demonstrations are in the range of (3,3)(-3,3)( - 3 , 3 ). On real data, the discriminator also required more frequent updates to provide more reliable signals (see Table 2).

### v2

@@ Appendix A Infinite Horizon Model Predictive Path Integral
In this section we present the full algorithm in detail, including MPPI as described in (Williams et al., 2017). Modifications to “conventional” MPPI for MPAIL are highlighted in blue. Where applicable, $(\mathbf{x})_{i}$ indicates the $i$th entry of $\mathbf{x}$ (in its first dimension, if $\mathbf{x}$ is a tensor).
@@
2:Number of trajectories to sample NN;
@@
14:for $j\leftarrow 0$ to J1J-1 do
15:  for $k\leftarrow 0$ to N1N-1 do $\triangleright$ Model rollouts and costing (parallelized)
@@
17:   for $t^{\prime}\leftarrow 0$ to H1H-1 do
18:     $a^{k}_{t^{\prime}}\sim\mathcal{N}((\mathbf{a}_{t})_{t^{\prime}},\Sigma)$ $\triangleright$ Sample action at predicted state
19:     $\tilde{s}^{k}_{t^{\prime}+1}\leftarrow f_{\psi}(\tilde{s}^{k}_{t^{\prime}},a^{k}_{t^{\prime}})$ $\triangleright$ Predict next state
20:     $c_{t^{\prime}}^{k}\leftarrow$ $c_{\theta}(s_{t^{\prime}}^{k},s_{t^{\prime}+1}^{k})$ $\triangleright$ Compute state-transition costs
@@
22:   $\mathcal{C}(\tau_{k})\leftarrow$ $\eta^{H}$$V_{\phi}(\tilde{s}_{H}^{k})$ + $\sum_{t^{\prime}=0}^{H-1}$ $\eta^{t^{\prime}}$$c_{t^{\prime}}^{k}$ $\triangleright$ Total trajectory cost
@@
26:  for $k\leftarrow 0$ to N1N-1 do $\triangleright$ Weight using exponential negative cost
@@
29:  for $t^{\prime}\leftarrow 0$ to H1H-1 do $\triangleright$ Optimal plan from weighted-average actions
@@
34:  $(\bm{\sigma}_{t})_{i}\leftarrow\sqrt{\sum_{k=0}^{N-1}w(\tau_{k})[((\mathbf{a}_{t}^{J})_{0})_{i}-(a^{k}_{0})_{i}]^{2}}$
@@
36:return $\mathbf{a}^{J}_{t}$, $\bm{\sigma}_{t}$
@@
9:  $s_{t}\sim\mathcal{T}(\cdot|s_{t-1},a_{t-1})$ $\triangleright$ Step and perceive environment
@@
12:   $a_{t}\sim\mathcal{N}((\mathbf{a}^{*}_{t})_{0},I\bm{\sigma}_{t})$
@@
16:  $r_{t}\leftarrow r_{\theta}(s_{t},s_{t+1})$ $\triangleright$ Reward from discriminator
@@ B.1 MPPI as a Policy
Let $f(s,s^{\prime})=c(s,s^{\prime})+\beta\mathbb{KL}(\pi(\cdot|s)||\overline{\pi}(\cdot|s))$. For either objective to be valid the cost and KL Divergence would have to be bounded. Thus, we may safely assume that $f$ is uniformly bounded $f||_{\infty}\leq K$. Let $\delta_{t}=\,\mathbb{E}_{s_{t},s_{t+1}\sim d^{t}}[f(s_{t},s_{t+1})]-\mathbb{E}_{s,s^{\prime}\sim d^{\pi}}[f(s,s^{\prime})]$ be the error between estimates of the objective.
@@
[…]
@@ B.2 MPAIL as an Adversarial Imitation Learning Algorithm
In this section, we integrate findings from Section B.1 with the AIL objective to theoretically validate MPAIL as an AIL algorithm. Specifically, we observe that at optimality, we recover the log expert-policy transition density ratio, which in turns yields a maximum entropy policy on state-transitions. We then discuss the identifiability limits imposed by observing only $(s,s^{\prime})$ rather than $(s,a,s^{\prime})$. Throughout this section we make use of the state-transition occupancy measure, defined as $\rho_{\pi}:\mathcal{S}\times\mathcal{S}\rightarrow\mathbb{R}$ where $\rho_{\pi}(s,s^{\prime})=\sum_{t=1}^{\infty}\gamma^{t}\mathcal{T}(S_{t+1}=s^{\prime},S_{t}=s\,|\,\pi)$ as in (Torabi et al., 2019a).
@@
Note that the optimal discriminator is achieved when $D^{\ast}(s,s^{\prime})=\frac{\rho_{E}(s,s^{\prime})}{\rho_{E}(s,s^{\prime})+\rho_{\pi}(s,s^{\prime})}$ as used in (Ghasemipour et al., 2020) and shown in (Goodfellow et al., 2014) Section 4 Proposition 1.
@@
Remark B.2.3. On Identifiability. A question naturally arises about the limitations that being state-only imposes. If state transitions are deterministic and invertible, observing $(s,s^{\prime})$ is the same as observing the unique action $a$ that caused it. Then $r(s,s^{\prime})=r(s,a)$ and by the 1-1 correspondence of policies with state-action occupancy measures (Ho and Ermon, 2016), the recovered policy becomes unique.
@@ C.1 Regularization
Spectral Normalization. As often found in GAN and AIL surveys (Orsini et al., 2021; Miyato et al., 2018), we corroborate that applying spectral normalization to the discriminator architecture appeared to have improved MPAIL training stability and performance. Application of spectral normalization to the value network did not appear to make a noticeable difference.
@@ C.2 Hyperparameters
Temperature Decay. As noted in Section 3, we found that an initial temperature with a gradual decay (down to a minimum) was helpful in preventing early and unrecoverable collapse. The intuition for this decision is similar to that of decaying policy noise injection in many popular RL frameworks (Lillicrap et al., 2019; Hansen et al., 2024), since the temperature is directly related to the variance of the optimized gaussian distribution. This component remains under investigation as its usage is not always necessary for meaningful convergence, but it is perhaps practically useful as it alleviates temperature tuning labor.
Value-to-Discriminator Update Ratio. Common to existing AIL (and GAN) implementations, MPAIL benefits from a balancing of generator and discriminator updates. Note that, like GANs, AIL tends to oscillate aggressively throughout training (Luo et al., 2024). As MPAIL does not enforce a constrained policy update each epoch (as TRPO does (Schulman et al., 2017a)), the policy is exposed more directly to the discriminator’s oscillations which can further hinder on-policy value estimation. A further converged value function is also theoretically more stationary from the perspective of $\pi_{\text{MPPI}}$ as infinite-horizon MPPI.
Markup. A notable quirk discovered during implementation is the relationship between costs and rewards. While the two concepts are generally regarded as dual (with negation), it is worthwhile noting that discounting is not closed under negation. Meaning, it is not correct to apply the same discount factor to the costs as they are done to the summation of rewards in the return $G_{t}$. Consider the reward with a discount applied $r_{1}=\gamma~r(s,s^{\prime})$ and the reward of a cost with a discount applied $r_{2}=-\gamma~c(s,s^{\prime})$. Observe that for $\gamma<1$, $r_{1}$ decreases while $r_{2}$ increases. Thus, when using costs, the $H$-step factor in the MPPI horizon, $\eta$, should not decrease over $t^{\prime}$. In fact, when applying $\eta<1$, we found that MPAIL does not meaningfully converge ever on the navigation task. (Geldenbott and Leung, 2024) names the usage of $\eta>1$ as a markup. In our case, we apply a similar empirical factor such that $\eta\coloneqq 1/\gamma>1$. While we suspect a more rigorous relationship between $\eta$ and $\gamma$, we leave its derivation for future work. However, we remark that a reward-only variant of MPPI which precludes these relationships is equally possible as done in (Hansen et al., 2022). Costs are maintained in this work due to wider familiarity in practice (Han et al., 2024a; Williams et al., 2017; Morgan et al., 2021; Finn et al., 2016).
@@ C.3 Computation
Figure 9: Comparing “Inference” Times for Navigation Task. Time taken to complete one episode of 100 timesteps with 64 parallel environments across varying horizon lengths and MPPI optimization iterations. PPO (in policy-based AIL) is used as implemented in the RSL library (Rudin et al., 2022). All training runs in this work are performed on an NVIDIA RTX 4090 GPU. Isaac Lab is chosen as our benchmarking and simulation environment due to its parallelization and robot learning extensions (Mittal et al., 2023; Han et al., 2025).
@@
(0): Linear(in_features=24, out_features=32,
bias=True)
@@
(2): Linear(in_features=32, out_features=32,
bias=True)
@@
(4): Linear(in_features=32, out_features=1,
bias=True)
@@
(0): Linear(in_features=12, out_features=32,
bias=True)
@@
(2): Linear(in_features=32, out_features=32,
bias=True)
@@
(4): Linear(in_features=32, out_features=1,
bias=True)
@@
Figure 10: PyTorch (Paszke et al., 2019) Model Architecture from Train Log. MPAIL readily admits other well-studied components of the model-based planning framework (e.g. sampling, dynamics) (Vlahov et al., 2025; 2024). This work focuses on costing from demonstration.
MPAIL is crucially implemented to be parallelized across environments in addition to trajectory optimization. In other words, in a single environment step, each parallel environment independently performs parallelized sampling, rollouts, and costing entirely on GPU without CPU multithreading. MPPI also allows for customize-able computational budget, similar to (Hansen et al., 2024) (see Figure 10). For the navigation and cartpole tasks, we find that online trajectory optimization implemented this way induces little impact on training times. In exchange, MPPI can be more space intensive due to model rollouts having space complexity of $\mathcal{O}(HN|\mathcal{S}|)$ per agent. On the navigation task benchmark settings, this is an additional 245 kB per agent or 15.7 MB in total for 64 environments. Figure 10 shows benchmarks on training times that demonstrate comparable times to PPO’s policy inference. Overall, training runs for Section 4.3 between MPAIL and GAIL on the navigation task are comparable at about 45 minutes each for 500 iterations.
@@ D.1 Real-Sim-Real Navigation
Setup Details. Before continuing with the discussion of our results, we provide further details about the setup of the experiment. The platform itself is an open-source MuSHR platform as detailed in (Srinivasa et al., 2023). Notably, the compute has been replaced with an NVIDIA Jetson Orin NX as mentioned in Section 4.2. Poses (position, orientation; $\begin{bmatrix}x&y&z&r&p&y\end{bmatrix}$) are provided by a motion-capture system at a rate of 20 Hz. Velocities are body-centric as estimated by onboard wheel encoders ($\mathbf{v}=v_{x}\mathbf{b}_{1}+v_{y}\mathbf{b}_{2}+v_{z}\mathbf{b}_{3}$, such that $\mathbf{b}_{1}$ points forward, $\mathbf{b}_{2}$ points left, and $\mathbf{b}_{3}$ completes the right-hand frame; basis vectors are rigidly attached to the vehicle (Han et al., 2024a)). Note that the vehicle is operated without slipping nor reversing such that $v_{y}\approx v_{z}\approx 0$ and $v_{x}>0$ (Han et al., 2024b). The recorded states used for the expert demonstration data is 240 timesteps long. Altogether, the data can be written as $s_{E}\in\{(x_{t},y_{t},z_{t},v_{x,t},v_{y,t},v_{z,t})\}_{t=1}^{240}$.
A remark: GAIL for this task is necessarily implemented with “asymmetry” between actor and reward. Since, the discriminator must receive as input expert observations $s_{E}$ while the agent is provided $(r,p,y)$ in addition to observations in $s_{E}$. In theory, there should be no conflict with the IRLfO (Equation 1) formulation as this remains a valid reward but on a subset of the state.
@@
To understand the role of partial observability in the experiment design, consider a simplified hand-designed cost using the partially observable expert data $c(s,s^{\prime}|s_{E},s_{E}^{\prime})$. A reference vector can be computed through the difference of positions between $s^{\prime}$ and $s$ then scaled by the demonstrated velocities: $c(s,s^{\prime}|s_{E},s_{E}^{\prime})\coloneqq\lVert^{\mathcal{I}}\mathbf{v}(s,s^{\prime})-{v_{x}}_{E}[(x_{E}^{\prime}-x_{E})\mathbf{e}_{1}+(y_{E}^{\prime}-y_{E})\mathbf{e}_{2}]\rVert_{2}$ where $\mathbf{e}_{i}$ are global basis vectors for global frame $\mathcal{I}$ and $^{\mathcal{I}}\mathbf{v}(s,s^{\prime})$ is the robot velocity in $\mathcal{I}$. Of course, this example assumes the ability to correctly choose the corresponding $(s_{E},s_{E}^{\prime})$ pair for input $(s,s^{\prime})$ out of the entirety of the expert dataset $d^{E}$. It should be clear that partial observability and state-transitions play critical roles in the recovery of this non-trivial cost function. This experiment presents a necessary challenge towards practical AIL (Orsini et al., 2021) and scalable Learning-from-Observation (LfO) (Torabi et al., 2019b).
@@
In both cases of GAIL and MPAIL, we find that the agents occasionally travel counter-clockwise (where the expert travels clockwise) during training, suggesting that $(s_{E},s_{E}^{\prime})$ appears close to $(s_{E}^{\prime},s_{E})$ through the discriminator. As the data is collected through real hardware, it is suspected that state estimation noise introduces blurring between states that are separated by only 50 ms. GAIL is otherwise known to perform poorly in the existence of multi-modal data (Li et al., 2017). This is further corroborated by its unstable performance on the navigation benchmark. And, to the best of our knowledge, similar Real-Sim-Real applications of AIL appear sparse if existent at all. Adjacent works which use real demonstration data but train in real include (Tsurumine and Matsubara, 2022; Sun et al., 2021). Even while training in real, GAIL’s performance drops signficantly ($90\%\rightarrow 20\%$) when presented with imperfect demonstrations for even straightforward tasks like reaching (Tsurumine and Matsubara, 2022; Sun et al., 2021). These observations might suggest why the GAIL discriminator is unable to learn meaningfully in simulation and produces a poor policy.